A jar contains 36 red marbles numbered 1 to 36 and 50 blue marbles numbered 1 to 50. A marble is drawn at random from the jar. Find the probability of the given event. Please enter reduced fractions.



(a) The marble is red. P(red)=





(b) The marble is odd-numbered. P(odd)=





(c) The marble is red or odd-numbered. P(red or odd) =





(d) The marble is blue or even-numbered. P(blue or even)


1
Expert's answer
2022-05-24T09:24:57-0400

(a)


"P(red)=\dfrac{36}{36+50}=\dfrac{18}{43}"

(b)


"P(odd)=\dfrac{36/2+50/2}{30+56}=\dfrac{1}{2}"

(c)


"P(red\ or\ odd)=\dfrac{36+50/2}{30+56}=\dfrac{61}{86}"

(d)


"P(blue\ or\ even)=\dfrac{36/2+50}{30+56}=\dfrac{34}{43}"